# Grounding Fallacies Misrepresenting Scientific Publications in Evidence

### v1

@@ F.3 Exhaustive Results on MissciPlus
Appendix G Details on Argument Reconstruction
G.1 LLM argument reconstruction
@@
We provide the LLM with all passages $S_{j}$ that were linked to at least one reasoning gap (or fallacy), and one passage $S^{0}$ that communicates the content based on which the claim was made. If multiple candidates for $S^{0}$ exist, we randomly select one. Experiments over Missci use the paraphrased $p_{0}$ instead of a randomly sampled passage $S^{0}$, and the publication context $s_{i}$ instead of the linked passages $S_{j}$. We run each experiment over three different seeds (1-3). For Llama3-8B we use a temperature of 0.3. For a fair assessment given the changed requirements for the LLM, we perform extensive prompt-search over all six prompts evaluated for Missci and an additional new prompts (cf. §I.3). During prompt search, we provide the complete fallacy information consisting of the definition, the logical form, and an example (cf. §A; Tables 7-8) within the instructions. All prompts task LLMs to output a ranked list of verbalized fallacious premises and the applied fallacy class. If possible, fallacies outside our inventory were converted; otherwise, they were removed. The results are listed in Table 22. We additionally report the average number of predicted fallacies per argument. This number excludes fallacies that the LLM hallucinated and are invalid (e.g., Contextomizer, Accident, Conflict of Interest, Fallacy of Conclusion) or that are outside of our fallacy inventory (e.g., Fear Mongering, Non Sequitur, Ad Hominem, False Consensus).
G.2 Llama3 Judge $\phi$
@@
We experiment with four different prompts (cf. §I.2) with the Llama3-8B-Instruct model as a binary classifier $\phi$ that determines whether two fallacious premises exhibit fallacious reasoning that bridges the identical gap. We use the human evaluation data provided by Glockner et al. (2024a) for training data. We discard all trivially invalid instances where the generated premise $\hat{\overline{p}}_{i}$ was almost identical to the claim or the paraphrased publication context $s_{i}$ by discarding all premise pairs if
@@ G.2.1 Prompting Strategies
G.2.2 Performance on Missci with only linked publication context
@@
Appendix H AFC applied on MissciPlus
H.1 Stance Predictions per Fallacy Class
@@
H.2 Binary Fallacy Detection by LLMs
@@
We adjusted our prompt (cf. §I.4) to allow LLMs to select no fallacy if they consider a claim correct and empirically evaluate three LLMs in how well they detect misinformation in MissciPlus or correct information in the 100 selected claims from HealthVer and CovidFact. Since true claims only come with one evidence passage, we also prompt the LLMs once only for misinformation using the concatenated relevant passages (cf. all passages in §6.2.1). Table 25 reports the ratio of claims identified as fallacious (the LLM did not specifically say that “no fallacy” exists and found at least one fallacy) across both datasets averaged over three seeds. We do not discern which fallacy was detected. Using the prompts to reconstruct fallacious arguments, the LLMs tend to find fallacies in all claims regardless of whether they are correct. We note that our claims are designed to correctly verbalize and identify all applied fallacies in MissciPlus. This follows previous works on fallacy detection (Da San Martino et al., 2019; Jin et al., 2022; Alhindi et al., 2022) that focused on which fallacy was applied rather than determining if a fallacy was applied. Future work may explore how to detect fallacies and correct claims better.
Appendix I Prompts
I.1 PRP Prompts
@@
I.2 Premise judge $\phi^{\mathrm{p+f}}$
@@
I.3 Argument reconstruction prompts
@@
We take the p1-p6 prompts from Missci as-is. For all-passages prompting we concatenate all passages (except for one $S^{0}_{j}$ passage), treating them as publication context in the Missci prompt, and treating the left-out $S^{0}_{j}$ as the accurate premise. The per-passage prompt templates for p7-p13 are shown in Figures 13-19., where we always replace [PASSAGE S0] with a randomly sampled passage $S_{j}^{0}$ and [PASSAGE Sj] with a passage $S_{j}$ linked to a fallacy. For all, we present the fallacy information consisting of the fallacy definition, logical form and example from literature identically to Missci prompts. The all-passage prompts are minimally edited from these prompts and have one dedicated field for all selected passages. All used prompts are provided within our repository.
I.4 LLM as AFC prompts
@@
The prompts used to directly predict the veracity of claims using LLMs are shown in Figures 20-22. We replace [EVIDENCE] with the concatenated evidence passages. The adapted fallacy generation prompt which allows to output that no fallacy exists, is shown in Figure 23.
Appendix J Reproducibility
All experiments with Llama2 or Llama3 used the instruction-tuned LLM and were performed on 80GB A100 GPUs. We always used the 70B and 8B models for Llama2 and Llama3, respectively. All experiments with LLMs are averaged over three seeds. The only exception is the experiments with GPT-4 Turbo in Table 3, which we only ran once due to computational costs. We used the API version 2023-10-01-preview for GPT-3.5 (model version: 0613) [^13] We used gpt-35-turbo-16k in the all passages prompting experiments and gpt-35-turbo in all other experiments, GPT-4 (model version: 0613) and GPT-4 Turbo (model version: 1106-Preview)
We used Grammarly [^14] https://app.grammarly.com/ in writing this paper.

### v2

@@ G.3 Exhaustive Results on MissciPlus
Appendix H Details on Argument Reconstruction
H.1 LLM argument reconstruction
@@
We provide the LLM with all passages $S_{j}$ that were linked to at least one reasoning gap (or fallacy), and one passage $S^{0}$ that communicates the content based on which the claim was made. If multiple candidates for $S^{0}$ exist, we randomly select one. Experiments over Missci use the paraphrased $p_{0}$ instead of a randomly sampled passage $S^{0}$, and the publication context $s_{i}$ instead of the linked passages $S_{j}$. We run each experiment over three different seeds (1-3). For Llama3-8B we use a temperature of 0.3. For a fair assessment given the changed requirements for the LLM, we perform extensive prompt-search over all six prompts evaluated for Missci and an additional new prompts (cf. §J.3). During prompt search, we provide the complete fallacy information consisting of the definition, the logical form, and an example (cf. §A; Tables 7-8) within the instructions. All prompts task LLMs to output a ranked list of verbalized fallacious premises and the applied fallacy class. If possible, fallacies outside our inventory were converted; otherwise, they were removed. The results are listed in Table 22. We additionally report the average number of predicted fallacies per argument. This number excludes fallacies that the LLM hallucinated and are invalid (e.g., Contextomizer, Accident, Conflict of Interest, Fallacy of Conclusion) or that are outside of our fallacy inventory (e.g., Fear Mongering, Non Sequitur, Ad Hominem, False Consensus).
H.2 Llama3 Judge $\phi$
@@
We experiment with four different prompts (cf. §J.2) with the Llama3-8B-Instruct model as a binary classifier $\phi$ that determines whether two fallacious premises exhibit fallacious reasoning that bridges the identical gap. We use the human evaluation data provided by Glockner et al. (2024a) for training data. We discard all trivially invalid instances where the generated premise $\hat{\overline{p}}_{i}$ was almost identical to the claim or the paraphrased publication context $s_{i}$ by discarding all premise pairs if
@@ H.2.1 Prompting Strategies
H.2.2 Performance on Missci with only linked publication context
@@
Appendix I AFC applied on MissciPlus
I.1 Stance Predictions per Fallacy Class
@@
I.2 Binary Fallacy Detection by LLMs
@@
We adjusted our prompt (cf. §J.4) to allow LLMs to select no fallacy if they consider a claim correct and empirically evaluate three LLMs in how well they detect misinformation in MissciPlus or correct information in the 100 selected claims from HealthVer and CovidFact. Since true claims only come with one evidence passage, we also prompt the LLMs once only for misinformation using the concatenated relevant passages (cf. all passages in §6.2). Table 25 reports the ratio of claims identified as fallacious (the LLM did not specifically say that “no fallacy” exists and found at least one fallacy) across both datasets averaged over three seeds. We do not discern which fallacy was detected. Using the prompts to reconstruct fallacious arguments, the LLMs tend to find fallacies in all claims regardless of whether they are correct. We note that our claims are designed to correctly verbalize and identify all applied fallacies in MissciPlus. This follows previous works on fallacy detection (Da San Martino et al., 2019; Jin et al., 2022; Alhindi et al., 2022) that focused on which fallacy was applied rather than determining if a fallacy was applied. Future work may explore how to detect fallacies and correct claims better.
I.3 LLMs may decline to provide a veracity prediction.
The percentages of the label classification from LLMs as fact-checking models using their parametric knowledge or RAG evidence in Table 6 may not sum up to 100% if the LLM declines to answer. We provide an extended Table that includes the percentage when the LLM declines to answer in Table 26.
Predicted as
LLM True False NEI No Answer True False NEI No Answer
Llama2 1.6 61.1 37.3 0.0 34.7 22.3 41.3 1.7
Know Llama3 0.0 86.9 2.4 10.7 20.0 43.3 14.3 22.3
(FC) GPT4 0.0 85.3 14.7 0.0 59.0 23.0 17.3 0.7
GPT3.5 0.8 71.0 28.2 0.0 46.7 17.3 35.7 0.3
Llama2 0.0 100.0 0.0 0.0 29.7 69.3 1.0 0.0
Know Llama3 8.3 88.9 2.4 0.4 68.7 26.0 3.0 2.3
(Ask) GPT4 3.6 68.3 27.8 0.4 49.7 6.7 36.3 7.3
GPT3.5 1.6 50.4 48.0 0.0 47.3 6.0 45.0 1.7
Llama2 23.8 61.5 12.7 2.0 58.7 29.7 10.7 1.0
RAG Llama3 44.4 53.2 2.4 0.0 80.3 16.3 3.3 0.0
GPT4 27.4 34.1 38.5 0.0 55.0 4.0 41.0 0.0
GPT3.5 38.9 31.7 29.0 0.4 78.0 5.3 16.0 0.7
Misinformation True claims
Appendix J Prompts
J.1 PRP Prompts
@@
J.2 Premise judge $\phi^{\mathrm{p+f}}$
@@
J.3 Argument reconstruction prompts
@@
We take the p1-p6 prompts from Missci as-is. For all-passages prompting we concatenate all passages (except for one $S^{0}_{j}$ passage), treating them as publication context in the Missci prompt, and treating the left-out $S^{0}_{j}$ as the accurate premise. The per-passage prompt templates for p7-p13 are shown in Figures 14-20., where we always replace [PASSAGE S0] with a randomly sampled passage $S_{j}^{0}$ and [PASSAGE Sj] with a passage $S_{j}$ linked to a fallacy. For all, we present the fallacy information consisting of the fallacy definition, logical form and example from literature identically to Missci prompts. The all-passage prompts are minimally edited from these prompts and have one dedicated field for all selected passages. All used prompts are provided within our repository.
J.4 LLM as AFC prompts
@@
The prompts used to directly predict the veracity of claims using LLMs are shown in Figures 21-23. We replace [EVIDENCE] with the concatenated evidence passages. The adapted fallacy generation prompt which allows to output that no fallacy exists, is shown in Figure 24.
Appendix K Reproducibility
All experiments with Llama2 or Llama3 used the instruction-tuned LLM and were performed on 80GB A100 GPUs. We always used the 70B and 8B models for Llama2 and Llama3, respectively. All experiments with LLMs are averaged over three seeds. The only exception is the experiments with GPT-4 Turbo in Table 4, which we only ran once due to computational costs. We used the API version 2023-10-01-preview for GPT-3.5 (model version: 0613) [^11] We used gpt-35-turbo-16k in the all passages prompting experiments and gpt-35-turbo in all other experiments, GPT-4 (model version: 0613) and GPT-4 Turbo (model version: 1106-Preview). We used Grammarly [^12] https://app.grammarly.com/ in writing this paper.